OpenStudy (anonymous):

The equation y = 5x2 + 6 is what kind of equation? A. Linear B. Rational C. Quadratic D. Exponential

OpenStudy (anonymous):

What do you have in mind?

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

What do you think the answer is?

OpenStudy (anonymous):

A

OpenStudy (uri):

I think it's c

OpenStudy (anonymous):

A linear function is in the y=mx+b form. This is not in that.

OpenStudy (anonymous):

Yes uri it is, but regan do you understand why it is?

OpenStudy (anonymous):

Its C quadratic equation :)

OpenStudy (anonymous):

Yes because it is squared, the square on the 5x makes it quadratic.

OpenStudy (anonymous):

Yes @KirbyLegs i understand

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Linear means when the power is one for example 2x+6=0 so here the power is 1 Quadratic means when the power is 2 for example 2x^2+6x+4=0 I hope I'm clear :)

OpenStudy (cwrw238):

the graph of a linear equation as you might escape is a straight line whereas a quadratic's is a curve